the sum of two numbers is 16 and the product is 63. what is the two numbers?
Answer by Maths68(1474)

Let numbers are x and y
x+y=16........(1)
xy=63
x=63y.........(2)
Put the value of x in (1)
63/y+y=16
Multiply by y both sides
63+y^2=16y
Rearrage above equation
y^2-16y+63=0
y^2-9y-7y+63=0
y(y-9)-7(y-9)=0
(y-7)(y-9)=0
y-7=0 or y-9=o
y=7 or y=9
Put the values of y in (1)
x=7
x+y=16
x+7=16
x=16-7
x=9
x+9=16
x=16-9
x=7
if one number is 7 other will be 9 or one is 9 other will be 7.
So numbers are 7 and 9